OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10)and (-4, 0)?

OpenStudy (anonymous):

choices 5/2, 2/5 -2/5 -5/2

OpenStudy (anonymous):

To find the slope when given two points you do \[(y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

where y sub 2 is the 2nd y-value and y sub 1 is the first y value (same for the x-values)

OpenStudy (y2o2):

\[(y2 - y1)\div(x2-x1)\]

OpenStudy (anonymous):

so we have 0 - (-10) = 0+10 = 10 then we have -4 - 0 = -4. So now we have 10 divide by -4 which simplifies to -5/2